OpenStudy (anonymous):

what is square root 2 over 2 exact answer

OpenStudy (anonymous):

you mean.. \[\frac{ \sqrt{2} }{ 2 }\] ?

OpenStudy (anonymous):

\[\frac{ \sqrt{2} }{ 2 }=\frac{{11} }{ c }\] trying to solve for c

OpenStudy (anonymous):

ill pm you

OpenStudy (anonymous):

\[\frac{ \sqrt{2} }{ 2 }=\frac{ 11\sqrt{3} }{ c }\] there it is! this is what im trying to figure out

OpenStudy (anonymous):

i know the answer is \[11\sqrt{6}\] but i dont know the steps how to get there

OpenStudy (anonymous):

cross multiply and you get: \[22\sqrt{3}=c \sqrt{2}\]

OpenStudy (anonymous):

The you divide \[\sqrt{2}\]

OpenStudy (anonymous):

on both sides and you get \[\frac{ 22\sqrt{3} }{ \sqrt{2} }\]

OpenStudy (anonymous):

but you can't have a square root on the bottom so then you multiply the square root 2 on the top and bottom

OpenStudy (anonymous):

so then you'll have \[\frac{ 22\sqrt{6} }{ 2 }\]

OpenStudy (anonymous):

you'll divide the two into the 22 and then you'll recieve your final answer \[11\sqrt{6}\]

OpenStudy (anonymous):

ahhh okay, that makes sense now, thank you, i've been stuck on this algebra

OpenStudy (anonymous):

glad i could help :)

OpenStudy (anonymous):

scotte12, I don't think this is algebra though.... ANYWAY, good that you get it.